Can you please help me
The width of a rectangle is 9 feet less than its length. The area of the rectangle is 90 ft^ . Find the length and width.
Please help me I wasn't here on the day they taught us this. Thanks

Answer by Alan3354(69443)

The width of a rectangle is 9 feet less than its length. The area of the rectangle is 90 ft^ . Find the length and width.
-----------
90 = W*L = W*(W + 9)
W^2 + 9W - 90 = 0
(W-6)*(W+15 = 0
W = 6
L = 15
============
Or try pairs of factors of 90
1*90
2*45
3*30
etc, until you find 2 that differ by 9